# Part-aware Shape Generation with Latent 3D Diffusion of Neural Voxel Fields

### v2

@@ 3.3 Part-aware Shape Decoder
where MHSA and MHCA represents the multi-head self-attention and cross-attention.
Incorporated with the part code, the refined voxel field $\overline{\mathbf{V}}^{\boldsymbol{f}}_{0}$ contains both the color and structural information. In this way, we can utilize the volume rendering to generate images and part maps of arbitrary views. Given a 3D position $\boldsymbol{x}$, we can query the point feature of the voxel field, and acquire the color and part emissions with MLP layers.
@@
where $\boldsymbol{p}\in\mathbb{R}^{K}$ is the part emission. In contrast to Eq. 3, we incorporate the part code into the MLP layers collectively to encode the structural information into individual points, thereby enhancing the generative intricacies. Subsequently, we derive the RGB image and semantic part map via ray casting following Eq. 2.
3.4 End-to-end Training with Gradient Skip
@@
where $\mathcal{L}_{rend}$ is the rendering loss. We ignore the term $\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}/\partial\mathbf{L}_{\boldsymbol{\phi}}$ to skip the gradient of decoder in the 3D autoencoder, resulting in a efficient formulation of gradient calculation:
@@
The total training objective is the summation of the diffusion loss and the rendering loss.
@@
4 Experiments
4.1 Experimental Setup
4.1.1 Dataset
We collect four categories of 3D objects from ShapeNet [5], including Chair, Table, Airplane and Car. Each category comprises 500 objects, with a split of 400 for training and 100 for testing purposes. Utilizing BlenderProc [6], we render 100 views of each object at a resolution of 800$\times$800 resolution. With the posed images, we use the public code base [41] to generate neural voxel fields.
4.1.2 Implementation Details
The resolution of the voxel field is 96$\times$96$\times$96, and the resolution of the latent code is 64$\times$ lower than the voxel field. We train the model for 100,000 iterations and we do not train the part-aware shape decoder for the first 20,000 iterations to warm up the latent 3D diffusion model. We set $K=4$ and $D=128$ for the part code. For each iteration, we sample 8192 rays to calculate the rendering loss. For the architecture of 3D UNet, we set the base channel to 96 and the channel multiplier is [1, 2, 4], resulting 144M parameters of the whole model.
4.1.3 Evaluation
We report the Frechet Inception Distance (FID) and Kernel Inception Distance (KID) to evaluate the quality of rendered images. We evaluate the two metrics at a resolution of 800$\times$800. In practice, we generate 200 objects and 10 views for each object to calculate metrics.
Table 1: Font sizes of headings. Table captions should always be positioned above the tables. Note the KID metrics are multiplied by 1000.
Method Chair Table Airplane Car
@@ 4.1.3 Evaluation
Part-NeRF 69.44 14.29 75.78 26.91 83.84 42.10 133.2 77.73
SSD-NeRF 49.15 8.56 - - - - 89.09 45.95
DiffRF 43.92 7.25 56.28 22.61 53.33 19.85 88.33 26.65
Ours 27.70 5.19 44.27 21.74 36.01 21.10 62.27 29.66
Table 2: Ablation study of the latent diffusion on Chair/Airplane. "LD" denotes "latent diffusion".
@@
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
with LD 27.70 5.19 36.01 21.10
without LD 47.19 8.08 57.65 22.05
Table 3: Ablation of part aware shape decoder on Chair/Airplane. "PASD" denotes "part-aware shape decoder".
@@
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
with PASD 27.70 5.19 36.01 21.10
without PASD 29.18 5.35 38.23 22.05
4.2 Shape Generation
Quantitative comparison against state of the art. In Tab. 1, we conduct a quantitative comparison between our proposed method and state-of-the-art approaches. Part-NeRF [42] employs multiple neural radiance fields to represent a single object, but the complex and redundant architecture increases the learning difficulty, ultimately yielding lower quality results. SSDNeRF [31] introduces a diffusion model for generating neural tri-plane fields. While the tri-plane representation is efficient, it lacks geometric detail compared to voxel representation. Additionally, our method employs a 3D UNet, which has superior feature extraction capabilities compared to the 2D UNet utilized in SSDNeRF, leading to enhanced performance. DiffRF [24] also utilizes the diffusion model in 3D space; however, it directly applies the 3D UNet to the voxel field. This implementation imposes significant computational demands, constraining the resolution of the voxel fields and leading to sub-optimal performances. We achieves the best FID metric on all classes, demonstrating the superiority of the proposed method.
Qualitative comparison against state of the art. As depicted in Fig. 3, we plot the visual generations of various state of the art methods. Leveraging the latent 3D diffusion model, we are able to generate high-resolution voxel fields, resulting in superior rendering quality. Similar to ours, Part-NeRF also produces part-aware rendering results. However, it employs an unsupervised approach that leads to the synthesis of meaningless structures. To ensure a fair comparison,we employ semantic part maps to supervise Part-NeRF, denoting this as Part-NeRF-S. Our approach still outperforms Part-NeRF-S in both texture and part quality. Both Quantitative and qualitative results indicate that the proposed method can generate accurate part-aware shapes and render high-quality images.
Figure 4: Visual improvements of different components. left: importance of latent diffusion; right: importance of part-aware shape decoder.
4.3 Ablation Studies
We ablate different components of the proposed method to show their importance. We conduct both quantitative and qualitative analyzes on the Chair data.
Importance of the latent 3D diffusion. In Fig. 2, we have demonstrated the importance of the 3D diffusion model. We further apply the diffusion process to the voxel field directly to observe the performance change. Due to the limitation of computational cost, we set the resolution of voxel fields to 32$\times$32$\times$32. Tab. 3 reports the performance comparison. Without the latent diffusion, the FID and KID drop a lot, which shows the importance of latent diffusion. Moreover, we visualize the qualitative results of the two different settings in Fig. 4, and the results reveal that the latent 3D diffusion brings high-quality rendering details.
Importance of the part-aware shape decoder. The part-aware shape decoder incorporates the part code into the neural voxel field, generating the structural shape with accurate rendering results. To demonstrate its importance, we replace it with two independent MLPs, rendering the color and part maps from the voxel field directly. Tab. 3 shows that the part-aware shape decoder can bring improvements in rendering quality. As depicted in Fig. 4, the part-aware shape decoder helps the model learn the accurate structural information.
@@ 4.3 Ablation Studies
4.4 Shape Interpolation
Following the 2D diffusion model [40], we can utilize two different voxel fields to generate intermediate shapes with spherical linear interpolation [36]. As shown in Fig. 6, our model can achieve relatively smooth interpolation between the two shapes. In contrast, the model without the part-aware shape decoder generates unreasonable contents, which further indicates the effectiveness of the part-aware shape decoder. With the part-aware shape decoder, the generated intermediate shapes consistently maintain a precise structure, suggesting that the decoder effectively learns to decompose the neural voxel fields with guidance of the part code.
4.5 Single Image to Shape
Single image to shape can be seen as a conditional generation task. In practice, we use Swin-S [21] as the image feature extractor to process the single image, and then concatenate the obtained features with the multi-scale features of the denoising UNet. As shown in Fig. 7, given a reference image as condition, our model can generate reasonable shapes with consistent geometry and accurate structure.
5 Conclusion
In this paper, we introduce a method for generating part-aware neural voxel fields using latent 3D diffusion models. The proposed method consists of a latent 3D diffusion model and a part-aware shape decoder. On one hand, the latent 3D diffusion model enhances the high-resolution of the neural voxel fields, thereby improving the geometric details and the quality of rendered images. On the other hand, the part-aware shape decoder integrates part codes into voxel fields to generate part-aware shapes with high-quality rendered images. We evaluate the effectiveness of our method on four classes by comparing it against state-of-the-art diffusion-based approaches, demonstrating its efficacy in generating precise part-aware shapes.
6 Limitation and Future Work
The proposed method achieves precise part-aware generation and high-quality rendering outcomes. However, acquiring 2D semantic part maps for supervision may pose challenges due to the difficulty in collecting labeled part maps. With the emergence of semantic big models like [18], we anticipate a significant reduction in this collection cost. Our goal is to implement pseudo-label based part-aware generation in future work.
7 More Training Details
Additional Training Loss. The main loss of training the network comprises two parts: diffusion loss $\mathcal{L}_{diff}$ and rendering loss $\mathcal{L}_{rend}$. Besides, we adopt two additional regularization loss functions. First, we conduct the total variation on the neural voxel fields to restrain the noise of generating:
$\mathcal{L}_{tv}=\sum_{dim\in[D]}{\sqrt{\Delta_{x}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)+\Delta_{y}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)+\Delta_{z}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)}},$ (12)
Sampling steps. Benefiting from fast sampling of the decoupled diffusion model, we can get satisfied results with only 30 steps. Tab. 4 reports the performance comparisons of different sampling steps for generating chairs.
Table 4: Performance comparisons of different sampling steps for generating chairs.
@@ 7 More Training Details
8 More Visual Results
We depict more comparisons with the supervision version of Part-NeRF (Part-NeRF-S) in Fig. 8. As a result, our method can generate higher-quality outcomes with more accurate structures. More generated samples can be found in Fig. 9 and Fig. 10.
Figure 8: Visual comparisons to Part-NeRF-S.
Figure 9: Generated samples of chair and airplane.
Figure 10: Generated samples of table and car.

### v3

@@ III-C Part-aware Shape Decoder
where MHSA and MHCA represent the multi-head self-attention and cross-attention.
Incorporated with the part code, the refined voxel field $\overline{\mathbf{V}}^{\boldsymbol{f}}_{0}$ contains both the color and structural information. In this way, we can utilize volume rendering to generate images and part maps of arbitrary views. Given a 3D position $\boldsymbol{x}$, we can query the point feature of the voxel field, and acquire the color emission and part class probability with MLP layers.
@@
where $\boldsymbol{p}\in\mathbb{R}^{K}$ is the part class probability. In contrast to Eq. 3, we incorporate the part code into the MLP layers collectively to encode the structural information into individual points, thereby enhancing the generative intricacies. Subsequently, we derive the RGB image and semantic part map via ray casting following Eq. 2.
III-D End-to-end Training with Gradient Skip
@@
where $\mathcal{L}_{rend}$ is the rendering loss. We ignore the term $\partial\hat{\mathbf{V}}^{\boldsymbol{f}}_{0}/\partial\mathbf{L}_{\boldsymbol{\phi}}$ to skip the gradient of the decoder in the 3D autoencoder, resulting in an efficient formulation of gradient calculation:
@@
The main training objective is the summation of the diffusion loss and the rendering loss.
@@
Except for the main training objective, we adopt two additional regularization loss functions. First, we conduct the total variation on the neural voxel fields to restrain the noise generating:
$\footnotesize\mathcal{L}_{tv}=\sum_{dim\in[D]}{\sqrt{\Delta_{x}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)+\Delta_{y}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)+\Delta_{z}^{2}(\hat{\mathbf{V}}^{\boldsymbol{f}}_{0},dim)}},$ (12)
IV Experiments
IV-A Experimental Setup
IV-A1 Dataset
We collect eight categories of 3D objects from ShapeNet [56], including Chair, Table, Airplane, Car, Lamp, Guitar, Laptop, and Pistol. For each category, we random sample 500 objects, with a split of 450 for training and 50 for testing purposes. Utilizing BlenderProc [57], we render 100 views of each object at a resolution of 800$\times$800 resolution. With the posed images, we use the public code base [48] to generate neural voxel fields. For text-guided shape generation, we get the text labels of Chair and Table from [58].
TABLE II: Quantitative comparisons to state-of-the-art methods on MMD and COV using Chamfer distance. Note the metrics are multiplied by 100.
Method Chair Table Airplane Car Guitar Lamp Laptop Pistol
@@ IV-A1 Dataset
Part-NeRF 1.17 60.4 1.26 56.4 0.48 39.6 0.64 23.8 0.26 31.7 1.57 59.4 0.86 31.7 0.34 48.5
EG3D 2.79 52.4 2.08 49.9 1.45 40.9 0.81 24.0 0.97 30.0 1.76 55.6 1.18 31.1 0.45 48.9
DiffRF 1.10 58.9 1.28 55.7 0.35 60.3 0.60 25.9 0.30 35.6 1.61 56.3 0.92 33.5 0.38 50.0
Ours 1.03 60.6 1.14 59.4 0.18 82.1 0.45 40.6 0.14 44.6 1.50 59.4 0.80 37.6 0.31 50.1
TABLE III: Quantitative comparisons to state-of-the-art methods on MMD and COV using Earth Mover’s distance. Note the metrics are multiplied by 100.
Method Chair Table Airplane Car Guitar Lamp Laptop Pistol
[…]
Part-NeRF 13.2 52.5 13.1 52.5 10.3 27.7 10.0 23.7 7.78 26.7 17.4 54.5 11.4 31.8 9.96 32.7
EG3D 20.2 38.5 18.5 47.6 9.39 42.0 14.2 22.1 10.5 30.3 21.9 43.9 16.0 31.9 9.98 34.1
DiffRF 13.5 50.8 13.4 54.3 9.95 30.3 9.41 28.8 7.01 30.5 17.8 53.9 13.4 32.0 9.52 36.1
Ours 12.3 53.1 12.4 57.4 6.99 51.5 8.78 36.6 6.49 36.5 16.2 55.5 11.0 37.8 9.34 36.7
Figure 4: Visual improvements of the part-aware shape decoder. Without the part-aware shape decoder, there are inconsistent rendering results in the yellow circles, which shows that part-aware information benefits texture learning.
Figure 5: Visual improvements of the latent diffusion.
IV-A2 Implementation Details
The resolution of the voxel field is 963, and the resolution of the latent code is 64$\times$ lower than the voxel field. We train the model for 100,000 iterations and we do not train the part-aware shape decoder for the first 20,000 iterations to warm up the latent 3D diffusion model. We set $K=4$ and $D=128$ for the part code. For each iteration, we sample 8192 rays of 5 random views to calculate the rendering loss. For the architecture of 3D UNet, we set the base channel to 96 and the channel multiplier is [1, 2, 4], resulting in 144M parameters of the whole model. We train the model in a single RTX 3090 GPU with batch size 2, taking up around 15, [^0] M GPU memory cost.
IV-A3 Conditioned Generation
We performed two types of modality for conditioned generation: image and text. For image condition, we use Swin-S [59] as the image feature extractor to process the single image, and then concatenate the obtained features with the multi-scale features of the denoising UNet. For text condition, we utilize CLIP [20] to extract the text embedding and correlate it with the multi-scale features of the denoising UNet via the cross-attention mechanism.
IV-A4 Evaluation
We report the Frechet Inception Distance (FID) and Kernel Inception Distance (KID) to evaluate the quality of rendered images. We evaluate the two metrics at a resolution of 800$\times$800. In practice, we generate 200 objects and 10 views for each object to calculate metrics. We also report the Coverage (COV) and the Minimum Matching Distance (MMD) [60] using Chamfer distance and Earth Mover’s distance. MMD measures how likely it is that a generated shape looks like a test shape. COV measures how many shape variations are covered by the generated shapes.
IV-B Shape Generation
Quantitative comparison against state of the art. In Tab. I, Tab. II and Tab. III, We have conducted a quantitative comparison between our proposed method and state-of-the-art approaches. Tab. I reports the metrics of generated image quality, while Tab. II and Tab. III report the geometry metrics calculated by Chamfer distance and Earth Mover’s distance. EG3D [13] presents a GAN-based framework that introduces a sophisticated hybrid explicit-implicit network architecture for synthesizing the triplane representation. Despite its lightweight nature, the triplane representation struggles to encompass intricate details, resulting in unsatisfied rendering quality and geometry. Part-NeRF [17] employs multiple neural radiance fields to represent a single object, but the complex and redundant architecture increases the learning difficulty, ultimately yielding lower-quality results. DiffRF [18] also utilizes the diffusion model in 3D space; however, it directly applies the 3D UNet to the voxel field. This implementation imposes significant computational demands, constraining the resolution of the voxel fields and leading to sub-optimal performances. Leveraging the 3D latent diffusion design enables the generation of 3D shapes with enhanced geometric details, leading to the attainment of the highest MMD and COV metrics across all classes. Additionally, the part-aware shape decoder facilitates the establishment of relationships between distinct parts, thereby improving the rendering of different components and resulting in the best FID metric across all classes.
Qualitative comparison against state of the art. As depicted in Fig. 3, we plot the visual generations of various state-of-the-art methods. Leveraging the latent 3D diffusion model, we can generate high-resolution voxel fields, resulting in superior rendering quality. Similar to ours, Part-NeRF also produces part-aware rendering results. However, it employs an unsupervised approach that leads to the synthesis of meaningless structures. To ensure a fair comparison, we employ semantic part maps to supervise Part-NeRF, denoting this as Part-NeRF-S. As shown in Fig. 3, benefiting from the design of latent diffusion and part-aware shape decoder, our approach still outperforms Part-NeRF-S in both texture and part quality. Both quantitative and qualitative results indicate that the proposed method can generate accurate part-aware shapes and render high-quality images.
TABLE IV: Ablation study of the latent diffusion on Chair/Airplane. ‘LD’ denotes ‘latent diffusion’. Here we use the Chamfer distance to calculate MMD and COV.
@@ IV-B Shape Generation
FID $\downarrow$ MMD $\downarrow$ COV $\uparrow$ FID $\downarrow$ MMD $\downarrow$ COV $\uparrow$
with LD 27.70 1.03 60.6 36.01 0.18 82.1
without LD 47.19 1.39 55.7 57.65 0.41 70.7
Figure 8: Text-guided shape generation. Given the specific descriptions, we can generate the corresponding shapes.
IV-C Ablation Studies
We ablate different components of the proposed method to show their importance.
TABLE V: Ablation of part aware shape decoder on Chair/Airplane. ‘PASD’ denotes ‘part-aware shape decoder’. Here we use the Chamfer distance to calculate MMD and COV.
@@ IV-C Ablation Studies
FID $\downarrow$ MMD $\downarrow$ COV $\uparrow$ FID $\downarrow$ MMD $\downarrow$ COV $\uparrow$
with PASD 27.70 1.03 60.6 36.01 0.18 82.1
without PASD 29.18 1.05 59.5 38.23 0.21 79.3
Importance of the latent 3D diffusion. Different from previous methods, this paper proposes the latent 3D diffusion that transfers the diffusion process into latent space, increasing the generative resolution significantly and preserving the high-quality generation. We further apply the diffusion process to the voxel field directly to observe the performance change. Due to the limitation of computational cost, we set the resolution of voxel fields to 32$\times$32$\times$32. Tab. IV reports the performance comparison. Without the latent diffusion, the FID drops a lot, and the geometry metric of MMD and COV drop 0.36 and 4.9, which shows the importance of latent diffusion. Moreover, we visualize the qualitative results of the two different settings in Fig. 5, and the results reveal that the latent 3D diffusion brings high-quality geometry and rendering details.
Importance of the part-aware shape decoder. The part-aware shape decoder incorporates the part code into the neural voxel field, generating the structural shape with accurate rendering results. To demonstrate its importance, we replace it with two independent MLPs, rendering the color and part maps from the voxel field directly. Tab. V shows that the part-aware shape decoder can bring more improvements in rendering quality. As depicted in Fig. 4, the part-aware shape decoder helps the model learn the accurate structural information, which allows for consistent rendering results of the local region.
Figure 10: Visual results of shape mixing. Given two part-ware shapes, we can directly compose the different parts to create new reasonable shapes.
@@
Importance of 3D UNet. We employ a 3D UNet as the network architecture to conduct the denoising process, and we have analyzed that the denoising network can be replaced by a 2D UNet possibly. However, we found that the 2D UNet failed to generate meaningful results as shown in Fig. 9. Compared to 3D UNet, the training loss curve of 2D UNet is not converged, culminating in a noisy generation. This observation indicates that it is necessary to utilize the 3D convolutional operator to extract features for the 3D voxel fields.
Effect of sampling steps. By leveraging rapid sampling from the decoupled diffusion model, we achieve satisfactory results in just 30 steps. Tab. VI presents a comparison of performance based on various sampling steps for chair generation. With an increase in the sampling step, performance shows a gradual improvement. Further increasing the sampling steps results in only a marginal improvement, suggesting that the generation quality is nearing saturation. It is observed that the sampling step has a greater impact on FID and KID metrics, highlighting its influence on image quality rather than geometry.
TABLE VI: Performance comparisons of different sampling steps for generating chairs.
@@
[…]
IV-D Shape Interpolation
Following the 2D diffusion model [50], we can utilize two different voxel fields to generate intermediate shapes with spherical linear interpolation [61]. As shown in Fig. 7, our model can achieve relatively smooth interpolation between the two shapes. In contrast, the model without the part-aware shape decoder generates unreasonable contents, which further indicates the effectiveness of the part-aware shape decoder. With the part-aware shape decoder, the generated intermediate shapes consistently maintain a precise structure, suggesting that the decoder effectively learns to decompose the neural voxel fields with the guidance of the part code.
IV-E Single Image to Shape
Single image to shape can be seen as a conditional generation task in which we use an image as conditional input and generate a shape related to the image. In practice, we extract the multi-scale image features via the Swin backbone network, and then concatenate the obtained features with the multi-scale features of the denoising UNet. As shown in Fig. 6, given a reference image as a condition, our model can generate reasonable shapes with consistent geometry and accurate structure.
IV-F Text-guided Shape Generation
Text-guided shape generation is a form of conditioned generative method, where the text is embedded into a feature to control the generation process. We perform text-guided shape generation on Chairs and Tables. As shown in Fig. 8, our model is capable of generating reasonable shapes based on the provided texts. In fact, the textual condition provides a relatively coarse level of control for generation, resulting in better geometric accuracy compared to texture detail.
IV-G Shape Mixing
Due to the part-aware nature of the generated shapes, it becomes feasible to combine parts from different shapes to create new designs. Notably, these shapes conform to a shared coordinate system, enabling the direct integration of disparate parts to create a cohesive and compact shape. Fig. 10 illustrates the visual results of shape mixing on four different object categories. Our approach excels in producing clear part shapes, facilitating the composition of diverse shapes.
V Conclusion
In this paper, we introduce a method for generating part-aware neural voxel fields using latent 3D diffusion models. The proposed method consists of a latent 3D diffusion model and a part-aware shape decoder. On one hand, the latent 3D diffusion model enhances the high resolution of the neural voxel fields, thereby improving the geometric details and the quality of rendered images. On the other hand, the part-aware shape decoder integrates part codes into voxel fields to generate part-aware shapes with high-quality rendered images. We evaluate the effectiveness of our method on four classes by comparing it against state-of-the-art diffusion-based approaches, demonstrating its efficacy in generating precise part-aware shapes.